A mass M is broken into two M and M-m these are separated by a distance r then find the ratio between M and M-m at which the gravitational force is maximum

Grade:12th pass

1 Answers

Aditya Gupta
5 years ago
We know that gravitational force is given by Gm1m2/r^2
Here m1=m and m2= M-m so force is proportional to m(M-m). For this to be maximum, we differentiate wrt m to obtain M-2m=0 or m=M/2
So ratio M/M-m= 2:1